# Anytime-valid simultaneous lower confidence bounds for the true discovery proportion

### v1

@@ 2 Notation and Related Literature
We consider the probability space $(\Omega,\mathcal{F},\mathcal{P})$, where $\Omega$ is a set of vectors of length $m$ and $\mathcal{P}$ is the set of all probability measures on $(\Omega,\mathcal{F})$. A probability distribution is denoted by $\mathbb{P}$ with expected value $\mathbb{E}_{\mathbb{P}}$. The probability measure under the true but unknown data generating process is denoted by $\mathbb{P}^{\ast}$, $\mathbb{P}^{\ast}\in\mathcal{P}$.
We aim to make inference on the set of hypotheses :-$\mathcal{H}\coloneq\{H_{1},\ldots,H_{m}\}$. We refer to $H_{i}$, $i\in\{1\ldots,m\}$, as elementary hypothesis. The considered hypotheses are composite hypotheses, that is, each elementary hypothesis $H_{i}$ corresponds to a set of probability measures $\mathcal{P}(i)\subset\mathcal{P}$. The hypothesis $H_{i}$ is true if and only if $\mathbb{P}^{\ast}\in\mathcal{P}(i)$.
Let $\mathcal{C}=2^{\{1,\ldots,m\}}\setminus\emptyset$ denote the collection of all nonempty subsets of $\{1,\ldots,m\}$. Intersection hypotheses are denoted by $H_{I}\coloneq\bigcap_{i\in I}H_{i}$, where $I\in\mathcal{C}$ is the set of indices corresponding to the elementary hypotheses in $H_{I}$. The size of set $I$ is denoted by $|I|$, with $|I|\geq 1$. Therefore $H_{I}$ might refer to an elementary or intersection hypothesis, while $H_{i}$ always denotes an elementary hypothesis. We denote by $\mathcal{P}(I)\coloneq\bigcap_{i\in I}\mathcal{P}(i)$ the set of all probability measures on $(\Omega,\mathcal{F})$ corresponding to $H_{I}$. The set $T\subseteq\{1,\ldots,m\}$ refers to the set of the indices of all true elementary hypotheses, which implies the assumption that $\mathbb{P}^{\ast}\in\mathcal{P}(T)$. Since $H_{T}$ includes all true elementary hypotheses, $\mathbb{P}^{\ast}\notin\mathcal{P}(i)$ for every $i\notin T$. Thus, an intersection hypothesis $H_{I}$ is true if and only if $I\subseteq T$.
@@
Lastly, let $\phi_{\alpha}$ denote a local level-$\alpha$ test which rejects an hypothesis iff $\phi_{\alpha}=1$ and which controls the type I error rate at level $\alpha$, i.e., $\mathbb{P}(\phi_{\alpha}(H_{I})=1)\leq\alpha$ for all $\mathbb{P}\in\mathcal{P}(I)$ and any $\alpha\in(0,1)$.
@@ 2.1 Simultaneous lower confidence bounds for the TDP
In this section, we provide more details about the construction of simultaneous lower confidence bounds for $\pi_{1}(R)$. Since $|R|$ is observable, $\tau(R)$ is the only unknown quantity in $\pi_{0}(R)$ and $\pi_{1}(R)$. Thus, to compute lower confidence bounds for $\pi_{1}(R)$, it is sufficient to compute upper confidence bounds for $\tau(R)$. Genovese and Wasserman (2006) and Goeman and Solari (2011) propose simultaneous upper confidence bounds for $\tau(R)$ based on closed testing. Goeman et al. (2021b) show that methods to compute confidence bounds for $\tau(R)$ are either based on closed testing or uniformly improved by closed testing. Closed testing is a class of multiple testing procedures proposed by Marcus et al. (1976). It can be used to determine a rejection set while controlling the family wise error rate, i.e., controlling the probability of making at least one false rejection. Denote by $\mathcal{X}_{\alpha}$ the set of indices of the hypotheses rejected by closed testing. Let $\mathcal{U}_{\alpha}$ be the set of indices corresponding to hypotheses that are rejected by a given local level-$\alpha$ test $\phi_{\alpha}$. The intersection hypothesis $H_{I}$ is rejected by closed testing based on $\phi_{\alpha}$, i.e., $I\in\mathcal{X}_{\alpha}$, if and only if for every $J\supseteq I$, $J\in\mathcal{U}_{\alpha}$. Goeman and Solari (2011) showed that
@@
$\mathbb{P}\left(\tau(R)\leq c_{\alpha}(R)\quad\forall R\in\mathcal{C}\right)\geq 1-\alpha,\qquad\text{for all }\mathbb{P}\in\mathcal{P}(T)\text{ and }\alpha\in(0,1).$ (2)
Equation (2) holds because $\mathbb{P}(T\notin\mathcal{U}_{\alpha})\geq 1-\alpha$ for all $\mathbb{P}\in\mathcal{P}(T)$ and therefore $\mathbb{P}(T_{R}\notin\mathcal{X}_{\alpha})\geq 1-\alpha$ for all $\mathbb{P}\in\mathcal{P}(T)$ and all $R\in\mathcal{C}$. In other words, the probability that $H_{T_{R}}$ is rejected by closed testing is less than $\alpha$, since this event requires that $H_{T}$ is rejected by $\phi_{\alpha}$, which has a probability of less than $\alpha$. Because the event $T\notin\mathcal{U}_{\alpha}$ does not depend on the discovery set $R$, the bounds are valid simultaneously for all $R\in\mathcal{C}$.
This approach requires testing up to $12^{m}-$ 1 hypotheses. When the local tests are based on p-values, Goeman et al. (2019) propose a shortcut which reduces the number of hypotheses that need to be tested explicitly.
@@ 2.2 E-variables and e-processes
An e-variable $E_{I}$ corresponding to hypothesis $H_{I}$ is a nonnegative random variable with $\mathbb{E}_{\mathbb{P}}[E_{I}]\leq 1$ for all $\mathbb{P}\in\mathcal{P}(I)$.
The realization of $E_{I}$ is called an e-value, we denote it by $e_{I}$. Because of Markov’s inequality, $\mathbb{P}(E_{I}\geq 1/\alpha)\leq\alpha$ for all $\mathbb{P}\in\mathcal{P}(I)$ and $\alpha\in(0,1)$. A local level-$\alpha$ test is thus given by $\phi_{\alpha}^{[E]}(H_{I})\coloneq\mathbb{I}\{e_{I}\geq 1/\alpha\}$, where $\mathbb{I}(\cdot)$ denotes the indicator function (Grünwald et al., 2024). It is possible to calibrate e-values based on p-values and vice-versa. Most notable, $P_{I}^{(E)}\coloneq\min(1/E_{I},1)$ is a valid, albeit conservative p-variable (Grünwald et al., 2024). However, the inverse of a p-variable does not need to be a valid e-variable. Data can be analyzed using both e-values and p-values, this has been investigated by Ignatiadis et al. (2024).
@@
There are many different families of e-variables. Most often, they are based on a likelihood ratio or Bayes factor (c.f. Grünwald et al., 2024; Wasserman et al., 2020). There are some guidelines on how to choose an e-variable in a specific setting. Mainly, the e-variable should grow as fast as possible if the evidence against the corresponding null hypothesis increases. Using these considerations, Grünwald et al. (2024) introduce the growth-rate optimal in the worst-case (GROW) criterion. This criterion measures how fast an e-variable can accumulate evidence against the null hypothesis under the worst-case alternative. It can be used to compare families of e-variables or specify e-variables.
@@
where $F$ is a symmetric e-merging function, such as the average. Denote by $\mathbf{d}^{R}_{F}$ the observed discovery e-vector. A discovery e-vector has the following interpretation: the number of true discoveries exceeds $j$ with a probability of at least $1-1/\mathbf{d}^{R}_{F}[j]$. These discovery e-vectors are the inverse of the confidence bounds by Goeman and Solari (2011). If $\mathbf{d}^{R}_{F}[j]\geq 1/\alpha$, then $c_{\alpha}(R)\leq|R|-j$. See Vovk and Wang (2023) for a detailed comparison between discovery e-vectors and the approach by Goeman and Solari (2011). Vovk and Wang (2024b) refine discovery e-vectors for independent and sequential e-values. A sequential e-variable is defined as a random variable with $\mathbb{E}_{\mathbb{P}}[E_{i}|E_{1},\ldots,E_{i-1}]\leq 1$ for all $\mathbb{P}\in\mathcal{P}(i)$. Sequential e-values can be utilized for temporal processes when the hypotheses are updated with time.
In our approach, we are interested in making inference on a set of hypotheses which is considered to be fixed at multiple observation time points, while allowing for optional stopping. This can be done using e-processes. Let $(\mathcal{F}_{n})_{n\geq 0}$ be a filtration, i.e., a sequence of sigma-algebras $\mathcal{F}_{0}\subset\mathcal{F}_{1}\subset\ldots\subset\mathcal{F}$.
@@
An e-process corresponding to $H_{I}$ is a nonnegative process $(E_{I}^{[n]})_{n\geq 0}$ adapted to some filtration $(\mathcal{F}_{n})_{n\geq 0}$ with $\mathbb{E}_{\mathbb{P}}[E^{[\nu]}_{I}]\leq 1$ for any $\mathcal{F}$-stopping time $\nu$ and for all $\mathbb{P}\in\mathcal{P}(I)$. (Ruf et al., 2023)
@@
If $H_{I}$ is true, the probability that a corresponding e-process ever exceeds a given threshold is bounded from above. Specifically, Ruf et al. (2023) showed that
$\mathbb{P}(\exists\nu\in\mathfrak{N}:E_{I}^{[n]}\geq\frac{1}{\alpha})\leq\alpha\quad\text{for every }\alpha\in[0,1]\text{ and every }\mathbb{P}\in\mathcal{P}(I).$ (3)
@@ 3 Anytime-valid simultaneous upper confidence bounds
We consider the filtered probability space $(\Omega,\mathcal{F},(\mathcal{F}_{n})_{n\in\mathbb{N}_{0}},\mathcal{P})$, with $\Omega$, $\mathcal{F}$ and $\mathcal{P}$ as in Section 2 and $(\mathcal{F}_{n})_{n\in\mathbb{N}_{0}}$ the natural filtration. That is, we assume that we sequentially observe data and at each observation time point we observe $m$ data points. We make no assumptions regarding the dependency structure of the data.
@@
Denote by $\mathcal{U}_{\alpha}^{[n]}$ the set of indices of the hypotheses rejected by $\phi_{\alpha}^{[E,n]}$ at time $n$. Furthermore, let $\mathcal{X}_{\alpha}^{[n]}$ be the set of indices corresponding to the hypotheses rejected by closed testing based on $\mathcal{U}_{\alpha}^{[n]}$. Utilizing $\phi_{\alpha}^{[E,n]}$ to determine $\mathcal{U}_{\alpha}^{[n]}$ requires valid e-processes for intersection hypotheses. It has been pointed out by Ramdas et al. (2023) that e-processes of elementary hypotheses can be merged, similarly to e-values. Lemma 1 demonstrates that $(E_{I}^{[n]})_{n\in\mathbb{N}_{0}}$ with $E_{I}^{[n]}=\sum_{i\in I}E_{i}^{[n]}/|I|$ is an e-process corresponding to the intersection hypothesis $H_{I}$.
@@
where $E_{i}^{[n]}$ is the e-process variable at time point $n$ corresponding to hypothesis $H_{i}$. For any stopping time $n\in\mathbb{N}_{0}$, $\mathbb{E}_{\mathbb{P}}[E_{I}^{[n]}]=\mathbb{E}_{\mathbb{P}}[\sum_{i\in I}E_{i}^{[\nu]}/|I|]=\sum_{i\in I}\mathbb{E}_{\mathbb{P}}[E_{i}^{[\nu]}/|I|\leq 1$ for all $\mathbb{P}\in\mathcal{P}(I)$ by definition of an e-process. Therefore, $(E_{I}^{[n]})_{n\in\mathbb{N}_{0}}$ is an e-process with respect to $(\mathcal{F}_{n})_{n\in\mathbb{N}_{0}}$ corresponding to $H_{I}$. ∎
@@
for all $\mathbb{P}\in\mathcal{P}(T)$ and $\alpha\in(0,1)$.
@@
Let $(E_{T}^{[n]})_{n\in\mathbb{N}_{0}}$ be the e-process corresponding to $H_{T}$. Because of Eq. (3), $\mathbb{P}(\exists n\in\mathbb{N}_{0}:E_{T}^{[n]}\geq 1/\alpha)\leq\alpha$ for every $\mathbb{P}\in\mathcal{P}(T)$ and every $\alpha\in[0,1]$ (Ruf et al., 2023). Thus, $\mathbb{P}(T\in\mathcal{U}_{\alpha}^{[n]})\leq\alpha$ for all $n\in\mathbb{N}_{0}$. Eq. (4) follows from the fact that $\tau(R)\leq c_{\alpha}^{[n]}(R)$ if $T_{R}\notin\mathcal{X}_{\alpha}^{[n]}$, which is the case if $T\notin\mathcal{U}_{\alpha}^{[n]}$. Since the event $T\notin\mathcal{U}_{\alpha}^{[n]}$ does not depend on the chosen discovery set $R$, the bounds are simultaneously valid for all $R\in\mathcal{C}$. ∎
@@
Computing $c_{\alpha}^{[n]}(R)$ at time point $n$ requires testing up to $12^{m}-$ 1 hypotheses, unless shortcuts are used. We propose such a shortcut when using the average as e-merging function and $\phi_{\alpha}^{[E,n]}$ to test the hypotheses. We assume that $R$ is fixed over time. The shortcut is based on the following lemma.
@@
Denote by $K_{h}^{[n]}\in\mathcal{C}_{R}$ the set of indices of the elementary hypotheses in a discovery set $R$ with the $h$ smallest e-process values at time point $n$, $n\in\mathbb{N}_{0}$. Denote by $\bar{K}_{h}^{[n]}\notin\mathcal{C}_{R}$ the set of indices corresponding to the elementary hypotheses not in $R$ with the $h$ smallest e-process values at time point $n$. Then, $K_{h}^{[n]}\notin\mathcal{X}_{\alpha}^{[n]}$ if
@@
[…]
@@
Remember that $K_{h}^{[n]}\notin\mathcal{X}_{\alpha}^{[n]}$ if there exists at least one $J\supseteq K_{h}^{[n]}$ such that $J\notin\mathcal{U}_{\alpha}^{[n]}$. Let $J\coloneq\{K_{h}^{[n]},\bar{K}_{k^{\ast}}^{[n]}\}$, where $\bar{K}_{k^{\ast}}^{[n]}=\emptyset$ is allowed. When using the local level-$\alpha$ e-test $\phi_{\alpha}^{[E,n]}$ and the average as an e-merging function, $J\notin\mathcal{U}_{\alpha}^{[n]}$ if
@@
Let $\mathcal{I}_{h}\subseteq\mathcal{C}_{R}$ be the set of indices of the discovery intersection hypotheses of size $h$, $h\in\{1,\ldots,|R|\}$. For every $h$, we are interested if there exists an hypothesis $H_{I_{h}}$ with $I_{h}\in\mathcal{I}_{h}$ and $I_{h}\notin\mathcal{X}_{\alpha}^{[n]}$ at time point $n$. This is true if and only if Eq. (5) holds. Furthermore, notice that the right side of Eq. (5) does not depend on $h$ and can be viewed as a discovery set specific threshold for testing any hypothesis $H_{I}$, $I\in\mathcal{C}_{R}$ at time $n$. Therefore, for a fixed discovery set $R$, we need to test at most $|R|$ hypotheses at time point $n$.

### v2

@@ 2 Notation and Related Literature
We consider the probability space $(\Omega,\mathcal{F},\mathcal{P})$, where $\Omega$ is a set of vectors of length $m$ and $\mathcal{P}$ is the set of all probability measures on $(\Omega,\mathcal{F})$. A probability distribution is denoted by $\mathbb{P}$ with corresponding expected value $\mathbb{E}_{\mathbb{P}}$. The probability measure under the true but unknown data generating process is denoted by $\mathbb{P}^{\ast}$, $\mathbb{P}^{\ast}\in\mathcal{P}$.
We aim to make inference on the set of hypotheses $\mathcal{H}\coloneq\{H_{1},\ldots,H_{m}\}$. We refer to $H_{i}$, $i\in\{1\ldots,m\}$, as elementary hypothesis. The considered hypotheses are composite hypotheses, that is, each elementary hypothesis $H_{i}$ corresponds to a set of probability measures $\mathcal{P}_{0}(i)\subset\mathcal{P}$. The hypothesis $H_{i}$ is true if and only if $\mathbb{P}^{\ast}\in\mathcal{P}_{0}(i)$.
Let $\mathcal{C}=2^{\{1,\ldots,m\}}\setminus\emptyset$ denote the collection of all nonempty subsets of $\{1,\ldots,m\}$. Intersection hypotheses are denoted by $H_{I}\coloneq\bigcap_{i\in I}H_{i}$, where $I\in\mathcal{C}$ is the set of indices corresponding to the elementary hypotheses in $H_{I}$. The size of set $I$ is denoted by $|I|$, with $|I|\geq 1$. Therefore, $H_{I}$ might refer to an elementary or intersection hypothesis, while $H_{i}$ always denotes an elementary hypothesis. We denote by $\mathcal{P}_{0}(I)\coloneq\bigcap_{i\in I}\mathcal{P}_{0}(i)$ the set of all probability measures on $(\Omega,\mathcal{F})$ corresponding to $H_{I}$. The set $T\subseteq\{1,\ldots,m\}$ refers to the index set of all true elementary hypotheses, which implies that $\mathbb{P}^{\ast}\in\mathcal{P}_{0}(T)$. Since $H_{T}$ includes all true elementary hypotheses, $\mathbb{P}^{\ast}\notin\mathcal{P}_{0}(i)$ for every $i\notin T$. Thus, an intersection hypothesis $H_{I}$ is true if and only if $I\subseteq T$.
@@
Lastly, let $\phi_{\alpha}$ denote a family of local level-$\alpha$ test which rejects hypothesis $H_{I}$ iff $\phi_{\alpha}(H_{I})=1$ and which controls the type I error rate at level $\alpha$, i.e., $\mathbb{P}(\phi_{\alpha}(H_{I})=1)\leq\alpha$ for all $\mathbb{P}\in\mathcal{P}_{0}(I)$ and any $\alpha\in(0,1)$.
@@ 2.1 Simultaneous lower confidence bounds for the TDP
In this section, we provide more details about the construction of simultaneous lower confidence bounds for $\pi_{1}(R)$. Since $|R|$ is observable, $\tau(R)$ is the only unknown quantity in $\pi_{0}(R)$ and $\pi_{1}(R)$. Thus, to compute lower confidence bounds for $\pi_{1}(R)$, it is sufficient to compute upper confidence bounds for $\tau(R)$. Genovese and Wasserman (2006) and Goeman and Solari (2011) propose simultaneous upper confidence bounds for $\tau(R)$ based on closed testing. Goeman et al. (2021b) show that methods to compute confidence bounds for $\tau(R)$ are either based on closed testing or uniformly improved by closed testing. Closed testing is a class of multiple testing procedures proposed by Marcus et al. (1976). It can be used to determine a rejection set while controlling the family wise error rate, i.e., controlling the probability of making at least one false rejection. Denote by $\mathcal{X}_{\alpha}$ the index set of the hypotheses rejected by closed testing. Let $\mathcal{U}_{\alpha}$ be the index set of the hypotheses that are rejected by a given local level-$\alpha$ test $\phi_{\alpha}$. The intersection hypothesis $H_{I}$ is rejected by closed testing based on $\phi_{\alpha}$, i.e., $I\in\mathcal{X}_{\alpha}$, if and only if for every $J\supseteq I$, $J\in\mathcal{U}_{\alpha}$. Goeman and Solari (2011) showed that
@@
$\mathbb{P}\left(\tau(R)\leq c_{\alpha}(R)\quad\forall R\in\mathcal{C}\right)\geq 1-\alpha,\qquad\text{for all }\mathbb{P}\in\mathcal{P}_{0}(T)\text{ and }\alpha\in(0,1).$ (2)
Equation (2) holds because $\mathbb{P}(T\notin\mathcal{U}_{\alpha})\geq 1-\alpha$ for all $\mathbb{P}\in\mathcal{P}_{0}(T)$ and therefore $\mathbb{P}(T_{R}\notin\mathcal{X}_{\alpha})\geq 1-\alpha$ for all $\mathbb{P}\in\mathcal{P}_{0}(T)$ and all $R\in\mathcal{C}$. In other words, the probability that $H_{T_{R}}$ is rejected by closed testing is less than $\alpha$, since this event requires that $H_{T}$ is rejected by $\phi_{\alpha}$, which has a probability of less than $\alpha$. Because the event $T\notin\mathcal{U}_{\alpha}$ does not depend on the discovery set $R$, the bounds are simultaneously valid for all $R\in\mathcal{C}$.
This approach requires testing up to $2^{m}-1$ hypotheses. When the local tests are based on p-values, Goeman et al. (2019) propose a shortcut which reduces the number of hypotheses that need to be tested explicitly.
@@ 2.2 E-variables and e-processes
An e-variable $E_{I}$ corresponding to hypothesis $H_{I}$ is a nonnegative random variable with $\mathbb{E}_{\mathbb{P}}[E_{I}]\leq 1$ for all $\mathbb{P}\in\mathcal{P}_{0}(I)$.
The realization of $E_{I}$ is called an e-value, we denote it by $e_{I}$. Because of Markov’s inequality, $\mathbb{P}(E_{I}\geq 1/\alpha)\leq\alpha$ for all $\mathbb{P}\in\mathcal{P}_{0}(I)$ and $\alpha\in(0,1)$. A local level-$\alpha$ test is thus given by $\phi_{\alpha}^{[E]}(H_{I})\coloneq\mathbb{I}\{e_{I}\geq 1/\alpha\}$, where $\mathbb{I}(\cdot)$ denotes the indicator function (Grünwald et al., 2024). It is possible to calibrate e-values based on p-values and vice-versa. Most notable, $P_{I}^{(E)}\coloneq\min(1/E_{I},1)$ is a valid, albeit conservative p-variable (Grünwald et al., 2024). However, the inverse of a p-variable does not need to be a valid e-variable. Data can be analyzed using both e-values and p-values, this has been investigated by Ignatiadis et al. (2024). There are many different families of e-variables. Most often, they are based on a likelihood ratio or Bayes factor (c.f. Grünwald et al., 2024; Wasserman et al., 2020). There are some guidelines on how to choose an e-variable in a specific setting. Mainly, the e-variable should grow as fast as possible if the evidence against the corresponding null hypothesis increases. Grünwald et al. (2024) introduce the growth-rate optimal in the worst-case (GROW) criterion. This criterion measures how fast an e-variable can accumulate evidence against the null hypothesis under the worst-case alternative. It can be used to compare families of e-variables or specify e-variables.
@@
where $F$ is a symmetric e-merging function, such as the average. Denote by $\mathbf{d}^{R}_{F}$ the observed discovery e-vector. A discovery e-vector has the following interpretation: the number of true discoveries exceeds $j$ with a probability of at least $1-1/\mathbf{d}^{R}_{F}[j]$. These discovery e-vectors are the inverse of the confidence bounds by Goeman and Solari (2011). If $\mathbf{d}^{R}_{F}[j]\geq 1/\alpha$, then $c_{\alpha}(R)\leq|R|-j$. See Vovk and Wang (2023) for a detailed comparison between discovery e-vectors and the approach by Goeman and Solari (2011). Vovk and Wang (2024b) refine discovery e-vectors for independent and sequential e-values. A sequential e-variable is defined as a random variable with $\mathbb{E}_{\mathbb{P}}[E_{i}|E_{1},\ldots,E_{i-1}]\leq 1$ for all $\mathbb{P}\in\mathcal{P}_{0}(i)$. Sequential e-values can be utilized for temporal processes when the hypotheses are updated with time.
In our approach, we are interested in making inference on a set of hypotheses, which is considered to be fixed, at multiple observation time points, while allowing for optional stopping. This can be done using e-processes. Let $(\mathcal{F}_{n})_{n\geq 0}$ be a filtration, i.e., a sequence of sigma-algebras $\mathcal{F}_{0}\subset\mathcal{F}_{1}\subset\ldots\subset\mathcal{F}$.
@@
An e-process corresponding to $H_{I}$ is a nonnegative process $(E_{I}^{[n]})_{n\geq 0}$ adapted to some filtration $(\mathcal{F}_{n})_{n\geq 0}$ with $\mathbb{E}_{\mathbb{P}}[E^{[\nu]}_{I}]\leq 1$ for any $\mathcal{F}$-stopping time $\nu$ and for all $\mathbb{P}\in\mathcal{P}_{0}(I)$. (Ruf et al., 2023)
@@
If $H_{I}$ is true, the probability that a corresponding e-process ever exceeds a given threshold is bounded from above. Specifically, Ruf et al. (2023) show that
$\mathbb{P}(\exists\nu\in\mathfrak{N}:E_{I}^{[\nu]}\geq\frac{1}{\alpha})\leq\alpha\quad\text{for every }\alpha\in[0,1]\text{ and every }\mathbb{P}\in\mathcal{P}_{0}(I).$ (3)
@@ 3 Anytime-valid simultaneous upper confidence bounds
We consider the filtered probability space $(\Omega,\mathcal{F},(\mathcal{F}_{n})_{n\in\mathbb{N}_{0}},\mathcal{P})$, with $\Omega$, $\mathcal{F}$ and $\mathcal{P}$ as in Section 2 and $(\mathcal{F}_{n})_{n\in\mathbb{N}_{0}}$ the natural filtration corresponding to the observable units. That is, we assume that we sequentially observe data and at each observation time point we observe $m$ data points. We make no assumptions regarding the dependency structure of the data.
@@
Denote by $\mathcal{U}_{\alpha}^{[n]}$ the index set of the hypotheses rejected by $\phi_{\alpha}^{[E,n]}$ at time $n$. Furthermore, let $\mathcal{X}_{\alpha}^{[n]}$ be the index set corresponding to the hypotheses rejected by closed testing based on $\mathcal{U}_{\alpha}^{[n]}$. Utilizing $\phi_{\alpha}^{[E,n]}$ to determine $\mathcal{U}_{\alpha}^{[n]}$ requires valid e-processes for intersection hypotheses. It has been pointed out by Ramdas et al. (2023) that e-processes of elementary hypotheses can be merged, similarly to e-values. Lemma 1 demonstrates that $(E_{I}^{[n]})_{n\in\mathbb{N}_{0}}$ with $E_{I}^{[n]}=\sum_{i\in I}E_{i}^{[n]}/|I|$ is an e-process corresponding to the intersection hypothesis $H_{I}$.
@@
where $E_{i}^{[n]}$ is the e-process variable at time point $n$ corresponding to hypothesis $H_{i}$. For any stopping time $n\in\mathbb{N}_{0}$, $\mathbb{E}_{\mathbb{P}}[E_{I}^{[n]}]=\mathbb{E}_{\mathbb{P}}[\sum_{i\in I}E_{i}^{[n]}/|I|]=\sum_{i\in I}\mathbb{E}_{\mathbb{P}}[E_{i}^{[n]}]/|I|\leq 1$ for all $\mathbb{P}\in\mathcal{P}_{0}(I)$ by definition of an e-process. Therefore, $(E_{I}^{[n]})_{n\in\mathbb{N}_{0}}$ is an e-process with respect to $(\mathcal{F}_{n})_{n\in\mathbb{N}_{0}}$ corresponding to $H_{I}$. ∎
@@
for all $\mathbb{P}\in\mathcal{P}_{0}(T)$ and $\alpha\in(0,1)$.
@@
Let $(E_{T}^{[n]})_{n\in\mathbb{N}_{0}}$ be the e-process corresponding to $H_{T}$. Because of Eq. (3), $\mathbb{P}(\exists n\in\mathbb{N}_{0}:E_{T}^{[n]}\geq 1/\alpha)\leq\alpha$ for every $\mathbb{P}\in\mathcal{P}_{0}(T)$ and every $\alpha\in[0,1]$ (Ruf et al., 2023). Thus, $\mathbb{P}(T\in\mathcal{U}_{\alpha}^{[n]})\leq\alpha$ for all $n\in\mathbb{N}_{0}$. Eq. (4) follows from the fact that $\tau(R)\leq c_{\alpha}^{[n]}(R)$ if $T_{R}\notin\mathcal{X}_{\alpha}^{[n]}$, which is the case if $T\notin\mathcal{U}_{\alpha}^{[n]}$. Since the event $T\notin\mathcal{U}_{\alpha}^{[n]}$ does not depend on the chosen discovery set $R$, the bounds are simultaneously valid for all $R\in\mathcal{C}$. ∎
@@
Computing $c_{\alpha}^{[n]}(R)$ at time point $n$ requires testing up to $2^{m}-1$ hypotheses, unless shortcuts are used. We propose such a shortcut when using the average as e-merging function and $\phi_{\alpha}^{[E,n]}$ to test the hypotheses. We assume that $R$ is fixed over time. The shortcut is based on the following lemma.
@@
Denote by $K_{h}^{[n]}\in\mathcal{C}_{R}$ the set of indices of the elementary hypotheses in a rejection set $R$ with the $h$ smallest e-process values at time point $n$, $n\in\mathbb{N}_{0}$. Denote by $\bar{K}_{h}^{[n]}\notin\mathcal{C}_{R}$ the set of indices corresponding to the elementary hypotheses not in $R$ with the $h$ smallest e-process values at time point $n$. Assume that $K_{h+1}^{[n]}\in\mathcal{X}_{\alpha}^{[n]}$. Then, $K_{h}^{[n]}\notin\mathcal{X}_{\alpha}^{[n]}$ if
@@
where $k^{\ast}=argmax\{k/\alpha-\sum_{j\in\bar{K}_{k}^{[n]}}e^{[n]}_{j}\}$, $0\leq k\leq m-|R|$, with $\sum_{j\in\bar{K}_{0}^{[n]}}e^{[n]}_{j}\coloneq 0$ by convention.
@@
Remember that $K_{h}^{[n]}\notin\mathcal{X}_{\alpha}^{[n]}$ if there exists at least one $J\supseteq K_{h}^{[n]}$ such that $J\notin\mathcal{U}_{\alpha}^{[n]}$. Let $J\coloneq\{K_{h}^{[n]},\bar{K}_{k^{\ast}}^{[n]}\}$, where $\bar{K}_{k^{\ast}}^{[n]}=\emptyset$ is allowed. Since $K_{h+1}^{[n]}\in\mathcal{X}_{\alpha}^{[n]}$, we do not consider any $J$ which includes additional rejected hypotheses, as we know that this hypothesis is rejected by closed testing. When using the local level-$\alpha$ e-test $\phi_{\alpha}^{[E,n]}$ and the average as an e-merging function, $J\notin\mathcal{U}_{\alpha}^{[n]}$ if
@@
Let $\mathcal{I}_{h}\subseteq\mathcal{C}_{R}$ be the index set of the rejected intersection hypotheses of size $h$, $h\in\{1,\ldots,|R|\}$. Lemma 2 allows us to sequentially test if there exists an hypothesis $H_{I_{h}}$ with $I_{h}\in\mathcal{I}_{h}$ for which Eq. (5) holds at time point $n$, with $h$ iteratively decreasing until $I_{h}\notin\mathcal{X}_{\alpha}^{[n]}$ and then $c_{\alpha}^{[n]}(R)=h$. Therefore, for a fixed rejection set $R$, we need to test at most $|R|$ hypotheses at time point $n$.
Furthermore, notice that the right side of Eq. (5) does not depend on $h$ and can be viewed as a rejection set specific threshold for testing any hypothesis $H_{I}$, $I\in\mathcal{C}_{R}$ at time $n$.